OpenStudy (anonymous):

CORRECT 4 MEDAL ^O^ Which of the following is the slope between the two points (2, 3) and (1, 4)? 1 1/3 -1 -1/3 <--- my answer

OpenStudy (jdoe0001):

\(\Large slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

4-3/1-2 = 1/1 = 1 ?

OpenStudy (anonymous):

or 4-3/1-2 = -1/3

jigglypuff314 (jigglypuff314):

4-3/1-2 = 1/-1 not 1/1 because 1-2 = -1

OpenStudy (anonymous):

@jigglypuff314 Thank you ! @jdoe0001 as well :D